Consider two markets: the market for motorcycles and the market for pancakes. The initial equilibrium for both markets is the same, the equilibrium price is $4.50, and the equilibrium quantity is 25.0. When the price is $6.75, the quantity supplied of motorcycles is 71.0 and the quantity supplied of pancakes is 105.0. For simplicity of analysis, the demand for both goods is the same.

Using the midpoint formula, calculate the elasticity of supply for pancakes. Please round to two decimal places.Consider two markets: the market for motorcycles and the market for pancakes. The initial equilibrium for both markets is the
Answer #1

The elasticity of supply for pancakes

= Q2-Q1/(Q2+Q1/2) / P2-P1/(P2+P1/2)

= 105-25/(105+25/2) / 6.75-4.50/(6.75+4.50/2)

= 80/65 / 2.25/5.625

= 1.2307/0.4

= 3.08

Supply in the market for motorcycle is - less elastic than supply in the market for pancakes

Elasticity in the mototcycle market = 71-25/(71+25/2) / 0.4

= 46/48 / 0.4

= 0.9583/0.4 = 2.39